Circular No. 7853

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002br IN MCG +03-27-61 M. Armstrong, Rolvenden, England, reports his discovery of an appar-

ent supernova (mag 17.1) on a 30-s unfiltered CCD image (limiting mag 18.5) taken on Mar. 9.970 UT with a 0.30-m Schmidt-Cassegrain reflector in the course of the U.K. Nova/Supernova Patrol. The new object is located at α = 10^{h}34^{m}08^{s}.92, δ = +19^{o}42′20′′.5, which is 5′′ north and 5′′ east of MCG +03-27-61. An image taken by T. Boles, Coddenham, England, on Mar. 13.843 shows the new object still at mag 17, and at position end figures 08^{s}.92, 21′′.3. The new object is not present on Armstrong’s images from 2000 Nov. 24, 2002 Jan. 17, Feb. 15, and Mar. 1, neither is it present on Palomar Sky Survey images taken on 1996 June 19 (limiting blue mag 22.5) and 1998 Jan. 2 (limiting red mag 20.8).

COMETS C/1998 A2, C/1998 A3, C/1999 M3, C/2000 B8 (SOHO) Further to IAUC 7850, D. Hammer reports his measurements for four

additional comets found by R. Kracht (1998, 1999 comets) and M. Meyer (C/2000 B8) on C2 coronagraph images posted at the SOHO website. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below; the 1998 and 1999 objects appear to belong to the Marsden group (cf. IAUC 7832), while the 2000 object is a member of the Meyer group.

Comet  1998 UT  α_{2000}  δ_{2000}  MPEC
C/1998 A2 Jan. 3.880  18^{h}54.^{m}5  −21^{o}54′  2002-E25
C/1998 A3  9.464  19 19.9  −21 14  2002-E25
Comet  1999 UT  α_{2000}  δ_{2000}  MPEC
C/1999 M3 June 30.396  6^{h}29.^{m}6  +23^{o}32′  2002-E18
Comet  2000 UT  α_{2000}  δ_{2000}  MPEC
C/2000 B8 Jan. 16.704  19^{h}56.^{m}5  −19^{o}59′  2002-E25

NOVA IN THE LARGE MAGELLANIC CLOUD 2002 Additional photometry (cf. IAUC 7847) by A. C. Gilmore at Mount

John Observatory: Mar. 11.514 UT, V = 12.04, U − B = −0.85, B − V = +0.14, V − R = +0.57, V − I = +0.85, airmass 1.43.

2002 March 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green